OpenStudy (anonymous):

A telephone pole casts a shadow of 1500cm. A girl who is 150cm tall is standing nearby. Her shadow is 125 cm long. How tall is the telephone pole?

OpenStudy (anonymous):

\[\frac{150}{125}=\frac{t}{1500}\]

OpenStudy (anonymous):

150*1500=125t

OpenStudy (anonymous):

t=225.000/125 t=1.800 telephone pole tall 1.800cm